A company’s costs, in millions of dollars, are given by the equation, C = x2 – 3x – 27, where x is the number of items sold, in thousands. What are the costs when 1,000 items are sold? 1,500 items?
Answer by richwmiller(17219)

C = x^2 – 3x – 27
For 1000 items substitute 1 for x
C = 1^2 – 3 – 27
C = 1– 30
-29 million dollars
For 1500 items substitute 1.5 for x
C = 1.5^2 – 3*1.5 – 27
C=2.25-4.5-27
c=2.25-31.5
c=-29.25 millions